The velocity of a moving vehicle is given by the equation "\ud835\udc63 = (2\ud835\udc61 + 3)^4". Use the Chain Rule to determine an equation for the acceleration when π‘Ž = 𝑑𝑣 𝑑𝑑 .


1
Expert's answer
2021-02-24T06:55:20-0500

Given: "v=(2t+3)^4"

Require to find "a=\frac{dv}{dt}" using Chain Rule.

Recollect the following (Chain Rule):

If "y=f(u)" and "u=g(x)" , then "\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}"

Let "u=(2t+3)"

Then "v=u^4,u=2t+3"

"\Rightarrow \frac{dv}{du}=4u^3,\frac{du}{dt}=2(1)+0=2"

Using Chain Rule, we get

"\frac{dv}{dt}=4u^3(2)=8u^3"

Substituting "u=2t+3," we get

"\frac{dv}{dt}=8(2t+3)^3"

Therefore the acceleration "a=\frac{dv}{dt}=8(2t+3)^3"